# When higher-order interactions enhance synchronization: the case of the Kuramoto model

### v1

@@ Abstract
Synchronization is a fundamental phenomenon in complex systems, observed across a wide range of natural and engineered contexts. The Kuramoto model provides a foundational framework for understanding synchronization among coupled oscillators, traditionally assuming pairwise interactions. However, many real-world systems exhibit group and many-body interactions, which can be effectively modeled through hypergraphs. Previous studies suggest that higher-order interactions shrink the attraction basin of the synchronous state, making it harder to reach and potentially impairing synchronization, despite enriching the dynamics. In this work, we show that this is not always the case. Through extensive numerical analysis of the higher-order Kuramoto model, we find that while strong higher-order interactions do generally work against synchronization, weak higher-order interactions can actually enhance it. This result is further corroborated by a cost-benefit analysis: under a constrained budget of both pairwise and higher-order interactions, a mixed allocation involving both consistently achieves higher synchronization than relying on either interaction type alone. These findings provide new insights into the role of higher-order interactions in shaping collective dynamics and point to design principles for optimizing synchronization in complex systems.
@@ Introduction
The key factor enabling this collective behavior are the interactions among the oscillators, which can be, e.g., mechanical (bridges, pendulum clocks), electrical (heart, power grids), visual (fireflies), or acoustic (frogs) Pikovsky et al. (2001); Strogatz (2003). Indeed, first Winfree understood that such interactions could be thought as a perturbation Winfree (1967), then Kuramoto Kuramoto (1975) obtained a simple model in which synchronization emerges as a phase transition triggered by the variation in the coupling strength: the celebrated Kuramoto model Strogatz (2000); Acebrón et al. (2005). In the classic Kuramoto model, two assumptions are taken which do not necessarily hold in applications. First, interactions are assumed to be pairwise, i.e., each one involving exactly two oscillators. Second, an all-to-all interaction topology is assumed, where each oscillator interacts with all others with the same strength.
@@
Over the past few years, it has been shown that higher-order interactions enrich system dynamics, with notable applications in the synchronization of chaotic oscillators Krawiecki (2014); Gambuzza et al. (2021); Gallo et al. (2022); Della Rossa et al. (2023), chimera states Kundu and Ghosh (2022); Muolo et al. (2024b), random walks Schaub et al. (2020); Carletti et al. (2020a), pattern formation Carletti et al. (2020b); Muolo et al. (2023), opinion dynamics Iacopini et al. (2019); Neuhäuser et al. (2020); DeVille (2021), and pinning control De Lellis et al. (2022); Xia and Xiang (2024); Muolo et al. (2025), to name a few. In particular, extensive literature shows that higher-order interactions enrich the dynamics of the Kuramoto model, generally making synchronization more difficult to achieve Tanaka and Aoyagi (2011); Bick et al. (2016); Skardal and Arenas (2019, 2020); Millán et al. (2020); Lucas et al. (2020); Adhikari et al. (2023); Skardal et al. (2023); León et al. (2024); Costa et al. (2024); Huh and Kim (2024); Wang et al. (2024); Zhang et al. (2024). In particular, Zhang et al. Zhang et al. (2024) showed that higher-order interactions make the attraction basin of the synchronous state smaller but more robust (deeper). These findings were further supported in Fariello and de Aguiar (2024), showing that the critical coupling for synchronization increases, facilitating desynchronization, and in von der Gracht et al. (2024), which demonstrated that once synchronization is achieved, it becomes harder to disrupt.
@@
These recent findings motivate our first research goal, which is to quantify the effects of higher-order interactions on the synchronous state for general hypergraph topologies. To this end, we studied Kuramoto oscillators coupled through $2$- and $3$-body interactions and conducted a detailed numerical analysis on various random hypergraphs. Our study confirms that, as expected, higher-order interactions enhance synchronization when the initial conditions are close to the synchronous state; however, when starting from incoherent states (i.e., far from synchronization), things become more interesting. While our simulations support the finding that higher-order interactions shrink the attraction basin of the synchronous state, they also reveal that weak higher-order interactions can, counterintuitively, enhance synchronization. As Zhang et al. Zhang et al. (2024) cleverly noted, the attraction basin becomes “deeper but smaller” with higher-order interactions. Our results seem to corroborate this observation and may be interpreted through the observation that the basin deepens before it begins to shrink.
@@
While the problem of optimally allocating links to promote specific collective behaviors, and in particular synchronization, has been extensively studied in complex networks, only a few studies have explored the role of higher-order interactions Lamata-Ot´ın et al. (2025). Much work focused on identifying metrics that promote or inhibit synchronization, showing, e.g., that factors contributing to good synchronizability are low eigenratio Pecora and Carroll (1998), high algebraic connectivity Coraggio et al. (2018), or high minimum density Coraggio et al. (2021, 2020). General network structural features also play a critical role: for example, small-world networks synchronize more effectively than random graphs Barahona and Pecora (2002). Notably, Donetti et al. Donetti et al. (2005, 2006) observed that optimally synchronizable networks tend to exhibit “entangled” homogeneous structures. Several approaches have also been proposed, from a network synthesis perspective, to optimally design synchronizable networks, both for oscillators ensembles Fazlyab et al. (2017); Lei et al. (2023) and for general dynamical systems Nishikawa and Motter (2006); Estrada et al. (2010); see also Coraggio et al. (2025) and references therein. Importantly, it has been shown that optimal network topologies can vary significantly when individual nodes dynamics are considered, highlighting a complex interplay between structure and dynamics, and, consequently, the coupling laws governing the interactions Gorochowski et al. (2010); Coraggio and di Bernardo (2024).
@@
Many of the mentioned studies focus on chaotic oscillators, whose synchronization is studied through the Master Stability Function, a linear stability analysis technique first proposed by Fujisaka and Yamada Fujisaka and Yamada (1983) and later extended to general networks by Pecora and Carroll Pecora and Carroll (1998), who also coined its current name. This approach, however, is not suited to coupled phase oscillators such as the Kuramoto model. On the other hand, the method exploited in Coraggio and di Bernardo (2024), can be also applied to phase oscillators and partly inspired the numerical approach we employed to answer this question. Our analysis shows that, when the total budget for interactions (pairwise and higher-order) is limited, synchronization is enhanced by a combination of both types of interactions, regardless of the relative cost of higher-order interactions within that budget.
The two main findings of this work—that (i) weak higher-order interactions enhance synchronization and that (ii) with a finite budget for connections, a combination of pairwise and higher-order interactions optimizes synchronization—are summarized in Figure 1.
@@
Figure 1: Main findings of this work: (1) While higher-order interactions generally hamper synchronization by making the attraction basin of the synchronous state “deeper but smaller”, weak higher-order interactions enhance synchronization. (2) Under a finite budget for interactions, regardless of the relative cost of higher-order interactions, the optimal configuration for synchronization always involves a combination of pairwise and higher-order interactions. The colors of the hyperedges give a pictorial representation of the coupling strength, namely red indicates a strong coupling while pink denotes a weaker one.
@@
In the next Section, we present the results of the two numerical studies on the higher-order Kuramoto model, focusing on the $3$-body interaction term for the $j$-th oscillator given by $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$. Results for a different $3$-body interaction, $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$ show qualitatively similar behavior and are included in the Supplementary Materials. In the Methods section, we provide rigorous definitions of the framework, detail the derivation of the higher-order Kuramoto model and the choice of interaction terms, and describe the numerical methods used.
@@ Results
To contextualize our results, we first briefly introduce the Kuramoto model with higher-order interactions that we considered in this study. We analyze a system of $N_{0}$ phase oscillators. The interactions are encoded by the adjacency matrix $A^{(1)}\in\{0,1\}^{N_{0}\times N_{0}}$ (i.e., the first order adjacency tensor) and the second order adjacency tensor $A^{(2)}\in\{0,1\}^{N_{0}\times N_{0}\times N_{0}}$, respectively. We assume undirected, unweighted, and unsigned interactions; i.e., connections are either present or absent. In particular, $A^{(1)}_{jk}=1$ if oscillators $j$ and $k$ are connected by a $2$-body interaction (a link), and 0 otherwise; $A^{(2)}_{jkl}=1$ if oscillators $j$, $k$ and $l$ are connected by a $3$-body interaction (a triangle), and 0 otherwise [^1] In the literature, a $2$-body interaction is equivalently referred to as a link, edge, pairwise interaction, $1$-hyperedge, or $1$-simplex. A $3$-body interaction is also called a triangle, $2$-hyperedge, or $2$-simplex. (see Methods for further details). The oscillators are assigned natural frequencies $\omega_{j}$, for $j\in\{1,...,N_{0}\}$, drawn randomly form a uniform distribution of a given interval. While this choice is arbitrary, we verified that whether the frequencies are drawn from a uniform or unimodal distribution does not affect the results (see also Methods), due to the relatively small size and heterogeneity of the hypergraphs. Hence, the higher-order Kuramoto model we consider is governed by the following equations, for $j\in\{1,\dots,N_{0}\}$:
$\dot{\vartheta_{j}}=\omega_{j}+\frac{K_{1}}{\langle d^{(1)}\rangle}\sum_{k=1}^{N_{0}}A_{jk}^{(1)}\sin(\vartheta_{k}-\vartheta_{j})+\frac{K_{2}}{2\langle d^{(2)}\rangle}\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j}),$ (1)
where $\langle d^{(1)}\rangle$ and $\langle d^{(2)}\rangle$ are the average $1$- and $2$-degree of $2$- and $3$-body interactions, respectively, and $K_{1}$ and $K_{2}$ are the coupling strengths of the $2$- and $3$-body interactions, respectively. The harmonic $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$ is not the only possible form of $3$-body coupling (see Methods); in the Supplementary Materials, we present analogous results for a different harmonic. The level of synchronization is measured by the Kuramoto order parameter $R\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{i\vartheta_{j}}\right\rvert$.
@@
To further validate the claim that the attraction basin of the synchronous state becomes “deeper but smaller” Zhang et al. (2024) as the strength of higher-order interactions increases, we conducted a systematic study of the higher-order Kuramoto model.
Let us now briefly illustrate the numerical procedure used to investigate the attraction basin of the synchronous state. More technical details can be found in the Methods section. First, we fix the total number of nodes ($N_{0}$), links ($N_{1}$) and triangles ($N_{2}$) and then generate $N_{\mathcal{H}}$ connected hypergraphs randomly allocating all links and triangles. We consider two connectivity conditions: $\mathcal{H}$-connectivity, if the projected network is connected, and $1$-connectivity, if the structure is already connected solely with the links. These two types of connectivity are illustrated in Figure 2, and their rigorous definitions are provided in the Methods section.
@@
We then ran the simulations varying the coupling strengths of pairwise and higher-order interactions, $K_{1}$ and $K_{2}$. For each pair, we performed $N_{\mathcal{H}}=300$ integrations over random hypergraphs, initial conditions and frequency distributions, averaging the resulting order parameter $R$ (see Methods for details). The results are shown in Figure 3.
@@
Figure 3: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order 33-body interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\vartheta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\vartheta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\vartheta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$: that value is never zero, a counterintuitive result. Top panels show the case of $\mathcal{H}$-connected hypergraphs, while bottom panels show the case of $1$-connected hypergraphs: we observe no significant differences between the two.
@@
The top panels show the case of $\mathcal{H}$-connected hypergraphs, while the bottom ones portray the case of $1$-connected hypergraphs. We see no significant differences, indicating that the observed effects arise from higher-order interactions, rather than connectivity differences. The left panels show the case in which the system starts close to the synchronized state. As expected from previous studies Zhang et al. (2024); von der Gracht et al. (2024), higher-order interactions enhance synchronization, as evidenced by the yellow regions (i.e., higher order parameter $R$) widening as $K_{2}$ increases. The right panels, instead, show the case of incoherent initial conditions. Again, as expected Zhang et al. (2024); Fariello and de Aguiar (2024); Muolo et al. (2025), we see that the stronger higher-order interactions ($K_{2}$) worsen synchronization (lower $R$). However, for small $K_{2}$ (weak higher-order interactions) $R$ initially increases before eventually declining. To better visualize this effect, we computed, for each $K_{1}$, the maximum value of $R$ in the case where the system starts from incoherence (right panels). Since weak higher-order interactions lead to higher values of $R$ also when starting from coherent initial conditions (left panels), we conclude that weak higher-order interactions enhance synchronization in general. We also tested $K_{1}$ and $K_{2}$ values up to $1.5$ (results omitted here for brevity), confirming the same trend. In the Supplementary Materials, we report a figure analogous to Figure 3, obtained assessing the Kuramoto-Daido order parameter $R_{2}\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{2i\vartheta_{j}}\right\rvert$, along with results for an alternative harmonic for the $3$-body interactions. All findings are consistent with those reported in the Main Text.
@@
Figure 4: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling in (1). The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\vartheta_{0}^{\mathrm{max}}]$$\vartheta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\vartheta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color). See Methods for further details. We observe that combining links and triangles generally yields higher synchronization than using either alone.

### v2

@@ Abstract
Synchronization is a fundamental phenomenon in complex systems, observed across a wide range of natural and engineered contexts. The Kuramoto model provides a foundational framework for understanding synchronization among coupled oscillators, traditionally assuming pairwise interactions. However, many real-world systems exhibit group and many-body interactions, which can be effectively modeled through hypergraphs. Previous studies suggest that higher-order interactions shrink the attraction basin of the synchronous state, making it harder to reach and potentially impairing synchronization, despite enriching the dynamics. In this work, we show that this is not always the case. Through a numerical study of higher-order Kuramoto models on random hypergraphs, we find that while strong higher-order interactions do generally work against synchronization, weak higher-order interactions can actually enhance it when combined with pairwise ones. This result is further corroborated by a cost-benefit analysis: under a constrained budget of both pairwise and higher-order interactions, a mixed allocation involving both consistently achieves higher synchronization than relying on either interaction type alone. These findings provide new insights into the role of higher-order interactions in shaping collective dynamics and point to design principles for optimizing synchronization in complex systems.
@@ Introduction
The key factor enabling this collective behavior are the interactions among the oscillators, which can be, e.g., mechanical (bridges, pendulum clocks), electrical (heart, power grids), visual (fireflies), or acoustic (frogs) Pikovsky et al. (2001); Strogatz (2003); Coraggio et al. (2025). Indeed, first Winfree understood that such interactions could be thought as a perturbation Winfree (1967), then Kuramoto Kuramoto (1975) obtained a simple model in which synchronization emerges as a phase transition triggered by the variation in the coupling strength: the celebrated Kuramoto model Strogatz (2000); Acebrón et al. (2005). In the classic Kuramoto model, two assumptions are taken which do not necessarily hold in applications. First, interactions are assumed to be pairwise, i.e., each one involving exactly two oscillators. Second, an all-to-all interaction topology is assumed, where each oscillator interacts with all others with the same strength.
@@
Over the past few years, it has been shown that higher-order interactions enrich system dynamics, with notable applications in the synchronization of chaotic oscillators Krawiecki (2014); Gambuzza et al. (2021); Gallo et al. (2022); Della Rossa et al. (2023), chimera states Kundu and Ghosh (2022); Muolo et al. (2024b), random walks Schaub et al. (2020); Carletti et al. (2020a), pattern formation Carletti et al. (2020b); Muolo et al. (2023), opinion dynamics Iacopini et al. (2019); Neuhäuser et al. (2020); DeVille (2021), and pinning control De Lellis et al. (2022); Xia and Xiang (2024); Muolo et al. (2025), to name a few. In particular, extensive literature shows that higher-order interactions enrich the dynamics of the Kuramoto model, generally making synchronization more difficult to achieve Tanaka and Aoyagi (2011); Bick et al. (2016); Skardal and Arenas (2019, 2020); Millán et al. (2020); Lucas et al. (2020); Adhikari et al. (2023); Skardal et al. (2023); León et al. (2024); Costa et al. (2024); Huh and Kim (2024); Wang et al. (2024); Zhang et al. (2024). In particular, Zhang et al. Zhang et al. (2024) showed that higher-order interactions make the attraction basin of the synchronous state smaller but more robust (deeper). These findings were further supported in Fariello and de Aguiar (2024), showing that the critical coupling for synchronization increases, facilitating desynchronization, and in von der Gracht et al. (2024); Wang et al. (2025a), where it was demonstrated that, once achieved, synchronization and twisted states becomes harder to disrupt due to higher-order interactions. Overall, while it is well known that stronger pairwise interactions enhance the degree of synchronization of phase oscillators in the purely pairwise case Kuramoto (1984); Acebrón et al. (2005); Rodrigues et al. (2016), the role of higher-order interactions remains far from fully understood.
@@
These recent findings motivate our first research goal, which is to quantify the effects of higher-order interactions on the synchronous state for more general hypergraph topologies. As in previous works Tanaka and Aoyagi (2011); Bick et al. (2016); Skardal and Arenas (2019, 2020); Millán et al. (2020); Lucas et al. (2020); Adhikari et al. (2023); Skardal et al. (2023); León et al. (2024); Costa et al. (2024); Huh and Kim (2024); Wang et al. (2024); Zhang et al. (2024); Fariello and de Aguiar (2024); von der Gracht et al. (2024); Wang et al. (2025a), our framework considers the addition of higher-order interactions to the pairwise setting. To this end, we studied Kuramoto oscillators coupled through $2$- and $3$-body interactions and conducted a numerical study on random hypergraphs of $10$ to $100$ nodes, a regime relevant for real-world applications (e.g., Menara et al. (2022); Simpson-Porco et al. (2013); Alderisio et al. (2017)). Our analysis confirms that, as expected, higher-order interactions generally enhance synchronization when the initial conditions are close to the synchronous state; however, when starting from incoherent states (i.e., far from synchronization), things become more interesting. While our simulations support the finding that higher-order interactions shrink the attraction basin of the synchronous state, they also reveal that weak higher-order interactions can enhance the degree of synchronization, when added to pairwise ones. As Zhang et al. Zhang et al. (2024) cleverly noted, the attraction basin becomes “deeper but smaller” with higher-order interactions. Our results seem to corroborate this observation and may be interpreted through the observation that the basin deepens before it begins to shrink.
@@
While the problem of optimally allocating links to promote specific collective behaviors, and in particular synchronization, has been extensively studied in complex networks, only a few studies have explored the role of higher-order interactions Lamata-Otín et al. (2025). Much work focused on identifying metrics that promote or inhibit synchronization, showing, e.g., that factors contributing to good synchronizability are low eigenratio Pecora and Carroll (1998), high algebraic connectivity Coraggio et al. (2018), or high minimum density Coraggio et al. (2021, 2020). General network structural features also play a critical role: for example, small-world networks synchronize more effectively than random graphs Barahona and Pecora (2002). Notably, Donetti et al. Donetti et al. (2005, 2006) observed that optimally synchronizable networks tend to exhibit “entangled” homogeneous structures, while Skardal et al. Skardal et al. (2014) found that matching between frequency heterogeneity and network heterogeneity enhances synchronization. Several approaches have also been proposed, from a network synthesis perspective, to optimally design synchronizable networks, both for oscillators ensembles Fazlyab et al. (2017); Lei et al. (2023) and for general dynamical systems Nishikawa and Motter (2006); Estrada et al. (2010); see also Coraggio et al. (2026) and references therein. Importantly, it has been shown that optimal network topologies can vary significantly when individual nodes dynamics are considered, highlighting a complex interplay between structure and dynamics, and, consequently, the coupling laws governing the interactions Gorochowski et al. (2010); Coraggio and di Bernardo (2024).
@@
Many of the mentioned studies focus on chaotic oscillators, whose synchronization is studied through the Master Stability Function, a linear stability analysis technique first proposed by Fujisaka and Yamada Fujisaka and Yamada (1983) and later extended to general networks by Pecora and Carroll Pecora and Carroll (1998), who also coined its current name. This approach, however, is not suited to coupled phase oscillators such as the Kuramoto model. On the other hand, the method exploited in Coraggio and di Bernardo (2024), can be also applied to phase oscillators and partly inspired the numerical approach we employed to answer this question. Our analysis on random hypergraphs shows that, when the total budget for interactions (pairwise and higher-order) is limited, synchronization is enhanced by a combination of both types of interactions, regardless of the relative cost of higher-order interactions within that budget.
The two main findings of this work—that (i) weak higher-order interactions enhance synchronization (i.e., increase the order parameter on average), when added to pairwise ones and that (ii) with a finite budget for connections, a combination of pairwise and higher-order interactions optimizes synchronization—were obtained via a numerical analysis of the higher-order Kuramoto model on random hypergraphs, and are summarized in Figure 1. These results are compatible with previous evidence for the all-to-all case, where it was shown that synchronization increases when higher-order interactions are added to pairwise ones Skardal and Arenas (2020); Kuehn and Bick (2021); Chen et al. (2025), and with those of two recent independent preprints, which have partially corroborated and complemented our settings by focusing on hyperrings Wang et al. (2025b); Skardal et al. (2025).
@@
Figure 1: a) Meaning of the order parameter $R$: high values correspond to synchronized states, whereas low values indicate incoherent dynamics. b) Pictorial representation of the main findings of this work. (1) While higher-order interactions generally hamper synchronization by making the attraction basin of the synchronous state “deeper but smaller”, adding weak higher-order interactions to pairwise-coupled networks enhances synchronization. (2) Under a finite budget for interactions, regardless of the relative cost of higher-order interactions, the optimal configuration for synchronization always involves a combination of pairwise and higher-order interactions. Red (resp. pink) hyperedges denote higher-order interactions with larger (resp. smaller) coupling strength. Yellow (resp. blue) nodes indicate highly (resp. weakly) synchronized states, in accordance with panel a).
@@
In the next section, we present the results of the two numerical studies on the higher-order Kuramoto model with $3$-body interactions. The functional form of such interactions is not unique; in the Main Text, we show results for one form, while results for a different one, show a qualitatively similar behavior and are included in the Supplementary Materials. In the Methods section, we provide rigorous definitions of the framework, detail the derivation of the higher-order Kuramoto model and the choice of interaction terms, and describe the numerical methods used.
@@ Results
To contextualize our results, we first briefly introduce the Kuramoto model with higher-order interactions that we considered in this study. We analyze a system of $N_{0}$ phase oscillators. Here, we present results for $N_{0}=10$; simulations for $N_{0}=100$—reported in the Supplementary Materials—exhibit the same qualitative behavior. The interactions are encoded by the adjacency matrix $A^{(1)}\in\{0,1\}^{N_{0}\times N_{0}}$ (i.e., the first order adjacency tensor) and the second order adjacency tensor $A^{(2)}\in\{0,1\}^{N_{0}\times N_{0}\times N_{0}}$, respectively. We assume undirected, unweighted, and unsigned interactions; i.e., connections are either present or absent.
In particular, $A^{(1)}_{jk}=1$ if oscillators $j$ and $k$ are connected by a $2$-body interaction (a link), and 0 otherwise; $A^{(2)}_{jkl}=1$ if oscillators $j$, $k$, and $l$ are connected by a $3$-body interaction (a triangle), and 0 otherwise [^1] In the literature, a $2$-body interaction is equivalently referred to as a link, edge, pairwise interaction, $1$-hyperedge, or $1$-simplex. A $3$-body interaction is also called a triangle, $2$-hyperedge, or $2$-simplex. (see Methods for further details). The oscillators are assigned natural frequencies $\omega_{j}$, for $j\in\{1,...,N_{0}\}$, drawn randomly from a uniform distribution of a given interval. Simulations with normally distributed frequencies are reported in the Supplementary Materials: we observed no qualitative difference between the two cases. Hence, the higher-order Kuramoto model we consider is governed by the following equations, for $j\in\{1,\dots,N_{0}\}$:
$\dot{\theta_{j}}=\omega_{j}+\frac{K_{1}}{\langle d^{(1)}\rangle}\sum_{k=1}^{N_{0}}A_{jk}^{(1)}\sin(\theta_{k}-\theta_{j})+\frac{K_{2}}{2\langle d^{(2)}\rangle}\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(\theta_{k}+\theta_{l}-2\theta_{j}),$ (1)
where $\langle d^{(1)}\rangle$ and $\langle d^{(2)}\rangle$ are the average $1$- and $2$-degree of $2$- and $3$-body interactions, respectively, and $K_{1}$ and $K_{2}$ are the coupling strengths of the $2$- and $3$-body interactions, respectively. Following Namura et al. (2025), we refer to the higher-order interaction term $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$ used in (1) as the (1,1,2)(1,1,-2) interaction. Results for a different $3$-body interaction, namely $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$, which we call the (2,1,1)(2,-1,-1) interaction, are reported in the Supplementary Materials, and show a qualitatively similar behavior to the simulations presented here. The level of synchronization is measured by the Kuramoto order parameter $R\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{i\theta_{j}}\right\rvert$.
@@
To further validate the claim that the attraction basin of the synchronous state becomes “deeper but smaller” Zhang et al. (2024) as the strength of higher-order interactions increases (in the presence of pairwise interactions), we conducted a systematic study of the higher-order Kuramoto model.
Specifically, we first fix the total number of nodes ($N_{0}$), links ($N_{1}$) and triangles ($N_{2}$) and then generate $N_{\mathcal{H}}$ connected hypergraphs randomly allocating all links and triangles. We consider two connectivity conditions: $\mathcal{H}$-connectivity, if the projected network is connected, and $1$-connectivity, if the structure is already connected solely with the links. These two types of connectivity are illustrated in Figure 2, and their rigorous definitions are provided in the Methods section. Then, we ran the simulations varying the coupling strengths of pairwise and higher-order interactions, $K_{1}$ and $K_{2}$. For each pair, we performed $N_{\mathcal{H}}=300$ integrations over random hypergraphs of $N_{0}=10$ nodes, initial conditions and frequency distributions, averaging the resulting order parameter $R$ (see Methods for details). Simulations for $N_{\mathcal{H}}=100$ integrations over random hypergraphs of $N_{0}=100$ nodes give analogous results, which are displayed in the Supplementary Materials.
@@
Figure 3: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order 33-body interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs, using the (1,1,2)(1,1,-2) interaction. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$: that value is never zero. Top panels show the case of $\mathcal{H}$-connected hypergraphs, while bottom panels show the case of $1$-connected hypergraphs: we observe no significant differences between the two.
Adding weak higher-order interactions enhances synchronization
@@
The top panels show the case of $\mathcal{H}$-connected hypergraphs, while the bottom ones portray the case of $1$-connected hypergraphs. We see no significant differences, indicating that the observed effects arise from higher-order interactions, rather than connectivity differences. The left panels show the case in which the system starts close to the synchronized state. As expected from previous studies Zhang et al. (2024); von der Gracht et al. (2024), higher-order interactions enhance synchronization, as evidenced by the yellow regions (i.e., higher values of the order parameter $R$) widening as $K_{2}$ increases. The right panels, instead, show the case of incoherent initial conditions. Again, as expected Zhang et al. (2024); Fariello and de Aguiar (2024); Muolo et al. (2025), we see that the stronger higher-order interactions ($K_{2}$) worsen synchronization (lower $R$). However, for small $K_{2}$ (weak higher-order interactions) $R$ initially increases before eventually declining. To better visualize this effect, we computed, for each $K_{1}$, the maximum value of $R$ in the case where the system starts from incoherence (right panels, black line).
Since weak higher-order interactions lead to higher values of $R$ also when starting from coherent initial conditions (left panels), we conclude that weak higher-order interactions enhance synchronization on average, when pairwise ones are already present.
We also tested $K_{1}$ and $K_{2}$ values up to $1.5$ (results omitted here for brevity), confirming the same trend.
In the Supplementary Materials, we report a figure analogous to Figure 3, obtained assessing the Kuramoto-Daido order parameter $R_{2}\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{2i\theta_{j}}\right\rvert$, along with results for the (2,1,1)(2,-1,-1) interaction for $3$-body interactions, normally distributed frequencies, and larger hypergraphs. All findings are consistent with those reported in the Main Text.
Emergence of distinct dynamical regimes
As shown in previous works Tanaka and Aoyagi (2011); Skardal and Arenas (2020); León et al. (2024); Battiston et al. (2025), higher-order interactions can induce bistability [^2] Coexistence of locally stable synchronized states and locally stable incoherent states. in the Kuramoto model. Consistently, Figure 3 shows a behavior compatible with bistability for low $K_{1}$ and high $K_{2}$: trajectories starting near the synchronous state remain synchronized, whereas incoherent initial conditions do not typically lead to synchronization.
To systematically explore the dependence of the system’s behavior on the coupling strengths $K_{1}$ and $K_{2}$, Figure 4 reports the distributions of the order parameter $R$ for representative parameter values across different realizations of hypergraphs, initial conditions, and frequency distributions (bottom panels), focusing on initially incoherent $1$-connected hypergraphs (Fig. 3, bottom-right panel). The top-left and top-right panels of Figure 4 show bifurcation diagrams of the average $R$ (blue line), together with its standard deviation (shaded area), obtained by varying one coupling strength while keeping the other fixed. An analogous analysis for Gaussian frequency distributions is provided in the Supplementary Materials.
For fixed $K_{1}=0.15$, large $K_{2}$ (pink upward triangle marker) yields a distribution of $R$ spanning both low and high values, which is compatible with bistability Suman and Jalan (2024). In contrast, for smaller $K_{2}$ (gray circle and pink downward triangle), despite a similar mean value of $R$, the distribution is more concentrated around the mean, indicating a more consistent behavior across realizations. For fixed $K_{2}=0.25$, increasing $K_{1}$ leads to distributions of $R$ that become increasingly concentrated near $1$, indicating a progressive stabilization of strongly coherent states across realizations.
Figure 4: System behavior with different pairwise ($K_{1}$) and higher-order ($K_{2}$) coupling strengths, for initially incoherent $1$-connected hypergraphs from Fig. 3. The top-center panel replicates the bottom right panel of Fig. 3 for comparison (color is the mean order parameter $R$). The top-left (resp. top-right) panel shows the mean value of $R$ (blue line) and its standard deviation (shaded area) as a function of the higher-order coupling strength $K_{2}$ (resp. $K_{1}$) while keeping the pairwise interaction strength $K_{1}$ (resp. $K_{2}$) fixed. The values of $K_{1}$, $K_{2}$ explored are also indicated in the top-center panel by a horizontal (resp. vertical) black line. Pink (resp. brown) triangles pointing upward/downward (resp. leftward/rightward) mark representative pairs $(K_{1},K_{2})$. The bottom panels display the distributions (integral normalized to 11) of $R$ across the $N_{\mathcal{H}}=300$ realizations of hypergraphs, initial conditions, and frequencies. For fixed $K_{1}=0.15$, a small $K_{2}$ (pink downward triangle) produces $R$ values relatively concentrated around the mean, whereas larger $K_{2}$ (pink upward triangle) leads to a broader distribution, spanning incoherent and strongly coherent regimes. For fixed $K_{2}=0.25$, increasing $K_{1}$ (brown leftward and upward triangles) causes the distribution to concentrate near $1$, indicating a consistently high level of synchronization.
@@
Figure 5: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the (1,1,2)(1,1,-2) interaction. $K_{1}$ and $K_{2}$ are the coupling strengths of $2$- and $3$-body interactions. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color). See Methods for further details. We observe that combining links and triangles generally yields higher synchronization than using either alone.